OpenStudy (anonymous):

A force of 20 newtons is exerted on a chair from its right and a force of 12 newtons is applied from its left. What is the net force acting on the chair and in what direction will the chair tend to move?

OpenStudy (anonymous):

From the following force diagram of the body:  we can write that the net force is: \[\Sigma F_x = 20 - 12 = 8(N)\] to the left (since the bigger force is applied from the right to the left). Thus, the body tends to move to the direction of the net force - to the left.